OpenStudy (anonymous):

Solve the following system of equations using substitution. u = 6 + t 36 = 2t + u A.t = 16, u = 22 B.t = 14, u = 20 C.t = 10, u = 16 D.t = 10, u = 4

OpenStudy (kirbykirby):

In the second equation, you can write: u=36-2t Thus equating the 1st and second equation: 6 + t = 36 - 2t t + 2t = 36 - 6 3t = 30 t = 10 Since t = 10, substitute this value into one of the equations, say the first one: u = 6 + t u = 6 + 10 u = 16